OpenStudy (anonymous):

The midpoint of CD is E(–1, 0). One endpoint is C(5, 2). What are the coordinates of the other endpoint?

OpenStudy (anonymous):

x+5 / 2 = -1

OpenStudy (amistre64):

you can draw a line to better see the distances involved

OpenStudy (amistre64):



OpenStudy (amistre64):

2+dy = 0 ; 0+dy = Y 5+dx = -1; -1+dx=X